If two dice are thrown what is the probability that the sum is
(1)greater than 8
(2)neither 7 nor 11
1
Expert's answer
2020-09-01T18:16:00-0400

(1)

Solution:

All possible variants of dices are "6\cdot6=36"

Suitable variants of dices are "10":

"(3,6)\\
(4,5),(4,6)\\
(5,4),(5,5),(5,6)\\
(6,3),(6,4),(6,5),(6,6)\\[0.2cm]"

Probability is "\dfrac{10}{36}=\dfrac{5}{18}".


Answer: 5/18


(2)

Solution:

All possible variants of dices are "6\cdot6=36"

Variants of dices which sum is 7 are "6":

"(1,6),(2,5),(3,4),(4,3),(5,2),(6,1)"

Variants of dices which sum is 11 are "2":

"(5,6),(6,5)"

Suitable variants of dices are all except 7 and 11:

"36-6-2=28\\[0.2cm]"

Probability is "\dfrac{28}{36}=\dfrac{7}{9}".


Answer: 7/9